### v1

@@ Abstract
We present the discovery of 11 new transiting brown dwarfs and low-mass M-dwarfs from NASA’s TESS mission: TOI-2844, TOI-3122, TOI-3577, TOI-3755, TOI-4462, TOI-4635, TOI-4737, TOI-4759, TOI-5240, TOI-5467, and TOI-5882. They consist of 5 brown dwarf companions and 6 very low mass stellar companions ranging in mass from 25 to 128 . We used a combination of photometric time-series, spectroscopic, and high resolution imaging follow-up as a part of the TESS Follow-up Observing Program (TFOP) in order to characterize each system. With over 50 transiting brown dwarfs confirmed, we now have a large enough sample to directly test different formation and evolutionary scenarios. We provide a renewed perspective on the transiting brown dwarf desert and its role in differentiating between planetary and stellar formation mechanisms. Our analysis of the eccentricity distribution for the transiting brown dwarf sample does not support previous claims of a transition between planetary and stellar formation at 42 . We also contribute a first look into the metallicity distribution of transiting companions in the range , showing that this too does not support a 42 transition. Finally, we also detect a significant lithium absorption feature in one of the brown dwarf hosts (TOI-5882) but determine that the host star is likely old based on rotation, kinematic, and photometric measurements. We therefore claim that TOI-5882 may be a candidate for planetary engulfment.
@@ 1 Introduction
Reframing our perspective on BDs into one motivated by formation and evolution has long been advocated for by some members of the BD community (Chabrier et al. 2014; Burrows 2014; Carmichael et al. 2021), where objects would be distinguished based on whether they form through a planet-like or a star-like formation mechanism. BDs forming like planets would undergo a core accretion pathway (Pollack et al. 1996), commonly referred to as a ”bottom-up” approach. The star-like BDs on the other hand would form via direct gravitational collapse, or ”top-down”, which can happen either within the circumstellar disk or at the core scale (Adams et al. 1989; Bate 2012; Kratter & Lodato 2016). Differentiating between these two formation pathways remains challenging, since it is unclear under what conditions each mechanism dominates, and whether there are any observable parameters that could distinguish them. Fortunately, in the era of TESS, we have begun to accumulate transiting BDs en masse, allowing us to pursue the question of BD formation from a different perspective. This budding population of transiting BDs is particularly enticing for studying BD formation because it provides a complementary, and in many cases, more complete understanding of the BD compared to the previously studied objects, which have primarily been discovered via direct imaging or radial velocity (RV) techniques. The transiting population serves as a complementary dataset to these other populations because transits provide a direct measurement of BD radii, a property which often can only be otherwise inferred with evolutionary models based on the observed spectrum and luminosity. This measurement is vital because BDs tend to contract with age, while also decreasing in size as mass increases (Baraffe et al. 2003; Saumon & Marley 2008; Burrows et al. 2001; Phillips et al. 2020). Thus there exists a degeneracy between mass, radius, and age for BDs making it difficult to test the substellar models with observed systems unless all three variables can be measured. These transiting systems provide direct, independent measurements on two of these degenerate parameters, and in cases where the host star’s age can be precisely determined, all three (e.g. Gillen et al. 2017; Nowak et al. 2017; David et al. 2019; Vowell et al. 2023).
@@
As this population of transiting brown dwarfs expanded in the era of space-based transit surveys, several new discoveries noted an ”oasis” forming in the desert (Carmichael et al. 2020; Šubjak et al. 2020; Henderson et al. 2024a) with new transiting systems beginning to populate the driest region of the desert noted by Ma & Ge (2014). With the new discoveries presented in this work, the transiting BD population now exceeds 50 systems, more than double the size of the population Ma & Ge (2014) had access to, opening the door for a reevaluation of brown dwarf desert from a new perspective. Hence, in this paper we present the discovery of 11 new transiting companions from NASA’s TESS mission. 6 of these systems are BD, with 3 lying within the Ma & Ge (2014) defined ”driest” region of the BD desert. We confirmed the remaining 6 non-BD, companions as very low-mass stars . In §2 of this manuscript, we present all the observations collected for each system in this work. §3 details our analysis of each system using EXOFASTv2 (Eastman et al. 2013, 2019). In §4 we provide a discussion on how these new systems fit into the population as a whole with a renewed perspective on the BD desert. We also discuss a detection of lithium in the host star of the BD companions presented here (TOI-5882). Finally, we present our conclusions in §5.
@@ 2 Observations
Right Ascension (h:m:s) | 07:20:48.778 | 15:07:48.989 | 21:48:23.012 | 04:38:59.373 | 1 | |
Declination (d:m:s) | 13:01:07.359 | -28:09:23.782 | 48:20:04.355 | 66:40:16.177 | 1 | |
@@
Right Ascension (h:m:s) | 18:18:40.777 | 02:14:31.26 | 06:53:38.51 | 06:23:44.23 | 1 | |
Declination (d:m:s) | 36:15:17.53 | 08:04:45.34 | -13:26:10.662 | -24:01:28.877 | 1 | |
@@
Right Ascension (h:m:s) | 19:32:20.11 | 06:17:34.49 | 20:47:33.291 | 1 | |
Declination (d:m:s) | 34:56:25.423 | 28:26:43.117 | 34:44:15.238 | 1 | |
@@ 2.1 TESS Photometry
The systems presented here were observed by TESS between Sectors with cadences ranging from 30-minutes in the prime mission to as low as 2-minutes in the extended mission. The TESS data were originally downloaded and reduced using both the TESS Science Processing Operations Center (SPOC) Pipeline (Jenkins et al. 2016) and the MIT Quick-Look Pipeline (QLP; Huang et al. 2020a, b; Kunimoto et al. 2021). The initial detection of a transit-like signal was discovered and vetted by the faint-star QLP search (Kunimoto et al. 2022) for 10 out the 12 systems presented in this paper, all but TOI-4462 and TOI-4635. These two TOIs were initially detected by the QLP and SPOC pipelines respectively, and then vetted by the TESS Science Office. The diagnostic tests described in Twicken et al. (2018) were used to evaluate whether the transit-like signal is indeed Keplerian. Upon passing, each system was designated as a TESS Object of Interest (TOI; Guerrero et al. 2021). While both the QLP and SPOC pipelines correct for contamination by known nearby stars, we choose to use the SPOC light curves with the shortest cadence in our analysis wherever possible for consistency. See Table 2 for full details on the sectors, cadence, and pipeline used for each source. It should be noted that all QLP light curves are processed from the full frame images. The TESS-SPOC (Caldwell et al. 2020) light curves are produced by the SPOC on a best-effort basis under the leadership of Doug Caldwell, the PI, and delivered as high-level science products to the Mikulski Archive for Space Telescopes (MAST) rather than as part of the official mission data products produced by the SPOC.
@@ 2.2 Ground-Based Time-Series Photometry
In order to confirm that the signal observed by TESS is on-target not originating from a nearby eclipsing binary that is blended with the target star, we gathered ground-based, time-series photometry of each system as the companion transited its host star. Since TESS has a relatively large pixel scale (21″ per pixel), the shallow eclipses we measured, which are consistent with roughly 1 , can be easily mimicked when a different, nearby eclipsing binary happens to fall on the same photometric aperture as the target star. The much deeper eclipses of the nearby eclipsing binary become diluted by the target star to mimic a much shallower event. Seeing-limited ground-based telescopes can have a much higher angular resolution than TESS typically 1-2″ and therefore can confirm that the signal is on-target, thereby ruling out nearby eclipsing binaries at all but the closest separations. It also has the benefit of observing in multiple wavelengths to confirm that the transit-like signal is achromatic. This is helpful because the eclipse depth of an eclipsing binary is nearly always wavelength dependent since the occulting body cannot be treated as a non-luminous sphere. We also note here, that while 6 of the companions presented here are low-mass M-dwarfs, and hence are eclipsing binaries themselves, they are so low in mass that we can still treat them as black spheres since they contribute negligibly to the overall flux of the system (Stevens et al. 2018). This process no only rules out the nearby eclipsing binary false positive, but also serves to refine the ephemerides of systems in which these data are able to extend the photometric baseline.
@@
All data sets, except for the observations of TOI-3577 from MuSCAT2, were reduced and their light curves extracted using AstroImageJ (AIJ; Collins et al. 2017). To do this, we use AIJ’s multi-aperture photometry tool using at least five similarly bright comparison stars. We use AIJ’s built-in transit fitting tool to assess the quality of the data and determine detrending parameters. Generally, we only detrend against the parameters that strongly correlate with the apparent brightness of the copmanion stars as they change over the course of the night. We also take care not to be too aggressive in our detrending, choosing to adopt detrending only when the Bayesian Information Criterion of AIJ’s transit-only fit significantly favors the detrended model. The detrending parameters used for each light curve can be found in Table 3. See §D in the appendix of Collins et al. (2017) for a detailed description of each detrending parameter. Finally, we normalized the data to the out-of-transit baseline and incorporated each light curve (with detrending) into our global fitting process (see §3).
@@ 2.3 TRES Spectroscopy
We obtained the spectroscopic measurements for each system in this sample via the Tillinghast Reflector Échelle Spectrograph (TRES) on the 1.5-meter Tillinghast Reflector telescope at the Fred Lawrence Whipple Observatory on Mt. Hopkins, Arizona. The TRES instrument is a fiber-fed, échelle spectrograph with a resolving power of 44,000. We reduced the spectra according to Buchhave et al. (2010) and analyzed each observation with the Stellar Parameter Classification (SPC) tool (Buchhave et al. 2012) in order to measure the metallicity, effective temperature, surface gravity, and projected rotational velocity of the star. We incorporated the average metallicity for each system into our analysis as a Gaussian prior in our global fits (see §3). We did not incorporate the effective temperature or surface gravity measurements from SPC as priors because these quantities are better constrained by the fit itself. This is due to the fact that EXOFASTv2 simultaneously models the spectral energy distribution, companion’s transit, and stellar evolutionary models (Eastman et al. 2023)
@@
Target | RV (m s-1) | (m s-1) | |
TOI-2844 | 2459528.8676 | -6696 | 1791 |
TOI-3122 | 2459651.9406 | 14841 | 145 |
TOI-3577 | 2459395.9184 | 52 | 70 |
TOI-3755 | 2459477.9662 | -8617 | 31 |
TOI-4462 | 2459468.6623 | 944 | 76 |
TOI-4635 | 2459556.7367 | 6170 | 19 |
TOI-4737 | 2459583.9325 | 2529 | 65 |
TOI-4759 | 2459623.7415 | -64 | 53 |
TOI-5240 | 2459681.9612 | 20204 | 375 |
TOI-5467 | 2459697.6358 | -237 | 352 |
TOI-5882 | 2459899.6240 | 108 | 55 |
@@ 2.4 High Resolution Imaging
The companion to TOI-4462 A was resolved by both SAI and PHARO at 0.4 separation and a position angle of 225 degrees. It is approximately 2.6 magnitudes dimmer in the and filters, too bright to neglect in our analysis (see §3). However, we remain confident that the Keplerian signals detected in both our photometry and spectroscopy can only be attributed to the brighter, primary star. The transits observed by TESS KeplerCam, and LCO-TEID show no evidence of chromaticity, and the spectral line profiles show no evidence of a secondary set of spectral lines that would produce an apparent RV shift. Hence we are confident that the signal we detect is due to an unresolved transiting companion around the brighter primary star TOI-4462 A.
@@ 3 Analysis
We analyzed each system using EXOFASTv2 444https://github.com/jdeast/EXOFASTv2 (Eastman et al. 2019), a publicly available exoplanet fitting suite. EXOFASTv2 is a Differential Evolution Markov Chain Monte Carlo (MCMC) code which globally fits both the star and the companion simultaneously, ensuring a self-consistent set of parameters for the entire system. In each fit, we generate a Spectral Energy Distribution (SED) model for the host star using MESA Isochrones and Stellar Tracks (MIST; Paxton et al. 2011, 2013) in order to fit the host star, while the companion is fit with a standard Keplerian model. Our SED model is fit to broadband archival photometry which we collected from Gaia Data Release 3 (DR3; Gaia Collaboration et al. 2023), 2MASS (Cutri et al. 2003; Skrutskie et al. 2006), and WISE (Wright et al. 2010; Cutri et al. 2012). The Keplerian model for the companion is fit to the TESS and ground based transits as well as the RV data from TRES. For a more detailed explanation of the modeling process, see Eastman et al. (2019).
@@
In addition to the priors described above, we also fit for a dilution term in each system to account for unresolved contaminants.To do this, we placed a prior of 0% 10% of the contamination ratio reported by the TESS Input Catalog (TIC; Stassun et al. 2018, 2019). While the QLP and SPOC light curves are both already corrected for known contaminants, we still chose to fit for a dilution term as a conservative assumption that the correction applied had a precision of at most 10%. We did this because the contamination ratio reported by the TIC is only an estimate that does not account for the actual point spread functions, as the CCD location and camera were unknown until after the launch of TESS. We also provided each fit with starting points on several parameters from the TIC. Specifically, we adopted the TIC-derived values for the host star’s mass, radius, and effective temperature as well as the companion’s orbital period, time of conjunction, and radius. We retrieved these values from the TESS mission catalog on ExoFOP555https://exofop.ipac.caltech.edu/tess/. We performed a preliminary fit with EXOFASTv2 on each system which included fitting a linear term to the radial velocities in order to account for a long-term drift due to unseen outer companions. In every case except for TOI-4737, this resulted in a slope consistent with zero within 1-sigma, and we subsequently fixed the slope to zero in all subsequent fits for these systems. For TOI-4737, we continued to fit for this long-term trend, and in the final iteration of these fits which we publish here, we found a slope of m s-1 day-1. Each system’s final fit was run to the adopted convergence criteria suggested by Eastman et al. (2019) of at least 1000 independent draws and a Gelman-Rubin statistic . See Table 6 for the priors used, and the median values determined from our analysis.
@@
(68.2% probability) | (31.8% probability) | (43.5% probability) | (56.5% probability) | ||
@@
Age | Age (Gyr) | ||||
EEP | Equal evolutionary phase | ||||
@@
Depth | TESSflux decrement at mid-transit | ||||

### v2

@@ Abstract
We present the discovery of 11 new transiting brown dwarfs and low-mass M-dwarfs from NASA’s TESS mission: TOI-2844, TOI-3122, TOI-3577, TOI-3755, TOI-4462, TOI-4635, TOI-4737, TOI-4759, TOI-5240, TOI-5467, and TOI-5882. They consist of 5 brown dwarf companions and 6 very low-mass stellar companions ranging in mass from 25 to 128 . We used a combination of photometric time-series, spectroscopic, and high-resolution imaging follow-up as a part of the TESS Follow-up Observing Program (TFOP) to characterize each system. With over 50 transiting brown dwarfs confirmed, we now have a large enough sample to directly test different formation and evolutionary scenarios. We provide a renewed perspective on the transiting brown dwarf desert and its role in differentiating between planetary and stellar formation mechanisms. Our analysis of the eccentricity distribution for the transiting brown dwarf sample does not support previous claims of a transition between planetary and stellar formation at 42 . We also contribute a first look into the metallicity distribution of transiting companions in the range , showing that this does not support a 42 transition too. Finally, we also detect a significant lithium absorption feature in one of the brown dwarf hosts (TOI-5882). However, we determine that the host star is likely old based on rotation, kinematic, and photometric measurements. We therefore claim that TOI-5882 may be a candidate for planetary engulfment.
@@ 1 Introduction
Reframing our perspective on BDs into one motivated by formation and evolution has long been advocated for by some members of the BD community (Chabrier et al. 2014; Burrows 2014; Carmichael et al. 2021), where objects would be distinguished based on whether they form through a planet-like or a star-like formation mechanism. BDs forming like planets would undergo a core accretion pathway (Pollack et al. 1996), commonly referred to as a ”bottom-up” approach. The star-like BDs on the other hand would form via direct gravitational collapse, or ”top-down”, which can happen either within the circumstellar disk or at the core scale (Adams et al. 1989; Bate 2012; Kratter & Lodato 2016). Differentiating between these two formation pathways remains challenging, since it is unclear under what conditions each mechanism dominates, and whether there are any observable parameters that could distinguish them. Fortunately, in the era of TESS, we have begun to accumulate transiting BDs en masse, allowing us to pursue the question of BD formation from a different perspective. This budding population of transiting BDs is particularly enticing for studying BD formation because it provides a complementary, and in many cases, more complete understanding of the BD compared to the previously studied objects, which have primarily been discovered via direct imaging or radial velocity (RV) techniques. The transiting population serves as a complementary dataset to these other populations because transits provide a model independent measurement of BD radii, a property which often can only be otherwise inferred with evolutionary models based on the observed spectrum and luminosity. This measurement is vital because BDs tend to contract with age, while also decreasing in size as mass increases (Baraffe et al. 2003; Saumon & Marley 2008; Burrows et al. 2001; Phillips et al. 2020). Thus there exists a degeneracy between mass, radius, and age for BDs making it difficult to test the substellar models with observed systems unless all three variables can be measured. These transiting systems provide direct, independent measurements on two of these degenerate parameters, and in cases where the host star’s age can be precisely determined, all three (e.g. Gillen et al. 2017; Nowak et al. 2017; David et al. 2019; Vowell et al. 2023).
@@
As this population of transiting brown dwarfs expanded in the era of space-based transit surveys, several new discoveries noted an ”oasis” forming in the desert (Carmichael et al. 2020; Šubjak et al. 2020; Henderson et al. 2024a) with new transiting systems beginning to populate the driest region of the desert noted by Ma & Ge (2014). With the new discoveries presented in this work, the transiting BD population now exceeds 50 systems, more than double the size of the population Ma & Ge (2014) had access to, opening the door for a reevaluation of brown dwarf desert from a new perspective. Hence, in this paper we present the discovery of 11 new transiting companions from NASA’s TESS mission. 6 of these systems are BDs, with 3 lying within the Ma & Ge (2014) defined ”driest” region of the BD desert. We confirmed the remaining 6 non-BD companions as very low-mass stars . In §2 of this manuscript, we present all the observations collected for each system in this work. §3 details our analysis of each system using EXOFASTv2 (Eastman et al. 2013, 2019). In §4 we provide a discussion on how these new systems fit into the population as a whole with a renewed perspective on the BD desert. We also discuss a detection of lithium in the host star of the BD companions presented here (TOI-5882). Finally, we present our conclusions in §5.
@@ 2 Observations
Right Ascension (h:m:s) | 07:20:48.78 | 15:07:48.99 | 21:48:23.01 | 04:38:59.37 | 1 | |
Declination (d:m:s) | 13:01:07.4 | -28:09:23.8 | 48:20:04.4 | 66:40:16.2 | 1 | |
@@
Right Ascension (h:m:s) | 18:18:40.78 | 02:14:31.26 | 06:53:38.51 | 06:23:44.23 | 1 | |
Declination (d:m:s) | 36:15:17.5 | 08:04:45.3 | -13:26:10.7 | -24:01:28.9 | 1 | |
@@
Right Ascension (h:m:s) | 19:32:20.11 | 06:17:34.49 | 20:47:33.29 | 1 | |
Declination (d:m:s) | 34:56:25.4 | 28:26:43.1 | 34:44:15.2 | 1 | |
@@ 2.1 TESS Photometry
The systems presented here were observed by TESS between Sectors with cadences ranging from 30-minutes in the prime mission to as low as 2-minutes in the extended mission. The TESS data were originally downloaded and reduced using both the TESS Science Processing Operations Center (SPOC) Pipeline (Jenkins et al. 2016) and the MIT Quick-Look Pipeline (QLP; Huang et al. 2020a, b; Kunimoto et al. 2021). The initial detection of a transit-like signal was discovered and vetted by the faint-star QLP search (Kunimoto et al. 2022) for 10 out the 12 systems presented in this paper. The remaining two, TOI-4462 and TOI-4635, were initially detected by the QLP and SPOC pipelines respectively, and then vetted by the TESS Science Office. The diagnostic tests described in Twicken et al. (2018) were used to evaluate whether the transit-like signal is indeed Keplerian. Upon passing, each system was designated as a TESS Object of Interest (TOI; Guerrero et al. 2021). While both the QLP and SPOC pipelines correct for contamination by known nearby stars, we choose to use the SPOC light curves with the shortest cadence in our analysis wherever possible for consistency. See Table 2 for full details on the sectors, cadence, and pipeline used for each source. It should be noted that all QLP light curves are processed from the full frame images. The TESS-SPOC (Caldwell et al. 2020) light curves are produced by the SPOC on a best-effort basis under the leadership of Doug Caldwell, the PI, and delivered as high-level science products to the Mikulski Archive for Space Telescopes (MAST) rather than as part of the official mission data products produced by the SPOC.
@@ 2.2 Ground-Based Time-Series Photometry
In order to confirm that the signal observed by TESS is on-target not originating from a nearby eclipsing binary that is blended with the target star, we gathered ground-based, time-series photometry of each system as the companion transited its host star. Since TESS has a relatively large pixel scale (21″ per pixel), the shallow eclipses we measured, which are consistent with roughly 1 , can be easily mimicked when a different, nearby eclipsing binary happens to fall on the same photometric aperture as the target star. The much deeper eclipses of the nearby eclipsing binary become diluted by the target star to mimic a much shallower event. Seeing-limited ground-based telescopes can have a much higher angular resolution than TESS typically 1-2″ and therefore can confirm that the signal is on-target, thereby ruling out nearby eclipsing binaries at all but the closest separations. It also has the benefit of observing in multiple wavelengths to confirm that the transit-like signal is achromatic. This is helpful because the eclipse depth of an eclipsing binary is nearly always wavelength dependent since the occulting body cannot be treated as a non-luminous sphere. We also note here, that while 6 of the companions presented here are low-mass M-dwarfs, and hence are eclipsing binaries themselves, they are so low in mass that we can still treat them as black spheres since they contribute negligibly to the overall flux of the system (Stevens et al. 2018). This process not only rules out the nearby eclipsing binary false positive, but also serves to refine the ephemerides of systems in which these data are able to extend the photometric baseline.
@@
All data sets, except for the observations of TOI-3577 from MuSCAT2, were reduced and their light curves extracted using AstroImageJ (AIJ; Collins et al. 2017). To do this, we use AIJ’s multi-aperture photometry tool using at least five similarly bright comparison stars. We use AIJ’s built-in transit fitting tool to assess the quality of the data and determine detrending parameters. Generally, we only detrend against the parameters that strongly correlate with the apparent brightness of the companion stars as they change over the course of the night. We choose to adopt detrending only when the Bayesian Information Criterion of AIJ’s transit-only fit significantly favors the detrended model. The detrending parameters used for each light curve can be found in Table 3. See §D in the appendix of Collins et al. (2017) for a detailed description of each detrending parameter. Finally, we normalized the data to the out-of-transit baseline and incorporated each light curve (with detrending) into our global fitting process (see §3).
@@ 2.3 Spectroscopy
We obtained spectroscopic measurements for each system in this sample via the Tillinghast Reflector Échelle Spectrograph (TRES) on the 1.5-meter Tillinghast Reflector telescope at the Fred Lawrence Whipple Observatory on Mt. Hopkins, Arizona. The TRES instrument is a fiber-fed, échelle spectrograph with a resolving power of 44,000. We reduced the spectra according to Buchhave et al. (2010) and analyzed each observation with the Stellar Parameter Classification (SPC) tool (Buchhave et al. 2012) in order to measure the metallicity, effective temperature, surface gravity, and projected rotational velocity of the star. We incorporated the average metallicity for each system into our analysis as a Gaussian prior in our global fits (see §3). We did not incorporate the effective temperature or surface gravity measurements from SPC as priors because these quantities are better constrained by the fit itself. This is due to the fact that EXOFASTv2 simultaneously models the spectral energy distribution, companion’s transit, and stellar evolutionary models (Eastman et al. 2023)
@@
2.3.1 SOPHIE Spectroscopy
We complement the TRES data with SOPHIE observations of TOI-5882. SOPHIE is a stabilized échelle spectrograph dedicated to high-precision radial-velocity measurements at the 1.93-m telescope of the Observatoire de Haute-Provence, France (Perruchot et al. 2008; Bouchy et al. 2009, 2013). We used its high-resolution mode (resolving power ) and the fast readout of its CCD.
Removing a few observations with low accuracy, we have a dataset of 21 SOPHIE measurements of TOI-5882 secured from December 2022 to July 2024. Exposure times ranged between 5 and 37 minutes, allowing signal-to-noise ratios between 14 and 40 to be reached per pixel at 550 nm.
The radial velocities were extracted with the SOPHIE pipeline, as presented by Bouchy et al. (2009) and refined by Heidari et al. (2024, 2025). That procedure includes corrections for bad pixels, cosmic rays, and charge transfer inefficiency of the CCD, as well as sky background and instrumental drifts. It derives cross correlation functions (CCF) from a numerical mask, then fit the CCFs by Gaussians to derive the radial velocities (Baranne et al. 1996; Pepe et al. 2002). One sample measurement is reported in Table 4.
Target | RV (m s-1) | (m s-1) | Spectrograph | |
TOI-2844 | 2459528.8676 | -6696 | 1791 | TRES |
TOI-3122 | 2459651.9406 | 14841 | 145 | TRES |
TOI-3577 | 2459395.9184 | 52 | 70 | TRES |
TOI-3755 | 2459477.9662 | -8617 | 31 | TRES |
TOI-4462 | 2459468.6623 | 944 | 76 | TRES |
TOI-4635 | 2459556.7367 | 6170 | 19 | TRES |
TOI-4737 | 2459583.9325 | 2529 | 65 | TRES |
TOI-4759 | 2459623.7415 | -64 | 53 | TRES |
TOI-5240 | 2459681.9612 | 20204 | 375 | TRES |
TOI-5467 | 2459697.6358 | -237 | 352 | TRES |
TOI-5882 | 2459899.6240 | 108 | 55 | TRES |
TOI-5882 | 2459930.2705 | -22361.2 | 12.4 | SOPHIE |
@@ 2.4 High Resolution Imaging
The companion to TOI-4462 A was resolved by both SAI and PHARO at 0.4 separation and a position angle of 225 degrees. It is approximately 2.6 magnitudes dimmer in the and filters, too bright to neglect in our analysis (see §3). However, we remain confident that the Keplerian signals detected in both our photometry and spectroscopy can only be attributed to the brighter, primary star. The transits observed by TESS KeplerCam, and LCO-TEID show no evidence of chromaticity, and the spectral line profiles show no evidence of a secondary set of spectral lines that would produce an apparent RV shift. Since TRES observed TOI-4462 with a 2.3 fiber (larger than the companion’s separation), the host star spectra were blended with light from the companion. However, this faint companion would only affect the measured RVs at the 10s of m/s level (Buchhave et al. 2011), significantly smaller than 10000 m/s semi-amplitude we measure for the TOI-4462 system. Hence we are confident that the signal we detect is due to an unresolved transiting companion around the brighter primary star TOI-4462 A.
@@ 3 Analysis
We analyzed each system using EXOFASTv2 444https://github.com/jdeast/EXOFASTv2 (Eastman et al. 2019), a publicly available exoplanet fitting suite. EXOFASTv2 is a Differential Evolution Markov Chain Monte Carlo (MCMC) code which globally fits both the star and the companion simultaneously, ensuring a self-consistent set of parameters for the entire system. In each fit, we generate a Spectral Energy Distribution (SED) model for the host star using MESA Isochrones and Stellar Tracks (MIST; Paxton et al. 2011, 2013) in order to fit the host star, while the companion is fit with a standard Keplerian model. Our SED model is fit to broadband archival photometry which we collected from Gaia Data Release 3 (DR3; Gaia Collaboration et al. 2023), 2MASS (Cutri et al. 2003; Skrutskie et al. 2006), and WISE (Wright et al. 2010; Cutri et al. 2012). The Keplerian model for the companion is fit to the TESS and ground based transits as well as the RV data from TRES and SOPHIE. For a more detailed explanation of the modeling process, see Eastman et al. (2019).
@@
In addition to the priors described above, we also fit for a dilution term in each system to account for unresolved contaminants.To do this, we placed a prior of 0% 10% of the contamination ratio reported by the TESS Input Catalog (TIC; Stassun et al. 2018, 2019). While the QLP and SPOC light curves are both already corrected for known contaminants, we still chose to fit for a dilution term as a conservative assumption that the correction applied had a precision of at most 10%. We did this because the contamination ratio reported by the TIC is only an estimate that does not account for the actual point spread functions, as the CCD location and camera were unknown until after the launch of TESS. We also provided each fit with starting points on several parameters from the TIC. Specifically, we adopted the TIC-derived values for the host star’s mass, radius, and effective temperature as well as the companion’s orbital period, time of conjunction, and radius. We retrieved these values from the TESS mission catalog on ExoFOP555https://exofop.ipac.caltech.edu/tess/ (NExScI 2022). We performed a preliminary fit with EXOFASTv2 on each system which included fitting a linear term to the radial velocities in order to account for a long-term drift due to unseen outer companions. In every case except for TOI-4737, this resulted in a slope consistent with zero within 1-sigma, and we subsequently fixed the slope to zero in all subsequent fits for these systems. For TOI-4737, we continued to fit for this long-term trend, and in the final iteration of these fits which we publish here, we found a slope of m s-1 day-1. Each system’s final fit was run to the adopted convergence criteria suggested by Eastman et al. (2019) of at least 1000 independent draws and a Gelman-Rubin statistic . See Table 6 for the priors used, and the median values determined from our analysis.
@@
(68.2% probability) | (31.8% probability) | (71.1% probability) | (28.9% probability) | ||
@@
Age (Gyr) | |||||
Equal evolutionary phase | |||||
@@
TESS flux decrement at mid-transit | |||||